OpenStudy (chillout):

Simple algebra questions: If 2^x*4^y = 32 and 3^x/9^y =3, what are the values for x and Y?

OpenStudy (jim766):

for the first one 2^x * 4^y = 32 think of the ways to mult 2 numbers and get 32 1 * 32, 2 * 16, 4 * 8 now pick the pair that you can use 8 * 4 = 32 2^3 * 4^1 = 32 x = 3, y = 1 make sense?

OpenStudy (chillout):

Indeed it does. But is there any other way besides trial and error?

OpenStudy (jim766):

probably, but I don't know it lol

OpenStudy (anonymous):

Trial and error would probably be your best bet:)

OpenStudy (chillout):

I mean, I can rewrite this as\[2^{x+2y}\ \ and\ \ 3^{x-2y}\] But I think it would get a bit tiring from here

OpenStudy (anonymous):

Yeahh it would. The trial and error method isn't 100% accurate but it's quite close

OpenStudy (jim766):

for the 2nd one, 9/3 what exponent can make each number its reciprocal

OpenStudy (jim766):

^ 9/3 = 3

OpenStudy (chillout):

Yes, I know, I was thinking in another method aside from this one, but thanks anyways. There will be some questions about it in a moment :)

OpenStudy (chillout):

Thanks for the help too, @Shawna.marie

OpenStudy (anonymous):

Nytime:)

OpenStudy (irishboy123):

there is no need to guess use your cleverly rewritten formula and also use logs. EG: (x+2y) ln(2) = ln32 (x-2y) ln(3) = ln3 x+2y = ln32/ln2 = 5 x-2y = 1 add them: 2x = 6: x=3, y = 1 check (3) - 2(1) = 1 and there you have it!!

OpenStudy (anonymous):

Omg that is like ninja status:)